OpenStudy (anonymous):

An object of mass 10 g is lifted a distance of 4 cm. How much work is done on the object? thanks!

OpenStudy (anonymous):

@AakashSudhakar

OpenStudy (anonymous):

@Jaynator495

OpenStudy (aakashsudhakar):

Step 1. Convert mass into kg. 10 g = 0.01 kg. Step 2. Convert distance into m. 4 cm = 0.04 m. Step 3. Work = F*x = mgx = (0.01kg)(9.8m/s^2)(0.04m) = 0.00392 N

OpenStudy (anonymous):

thanks !! got one more question for you @AakashSudhakar ..i wil tag you